OpenStudy (anonymous):

f(x)=x+1 g(x)=x^2-6x+8 find (f/g)(x) find (f/g)(1)

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

2/15

OpenStudy (anonymous):

\[\frac{f}{g}(x)=\frac{f(x)}{g(x)}=\frac{x+1}{x^2-6x+8}\]

OpenStudy (anonymous):

\[\frac{f}{g}(1)=\frac{1+1}{1^2-6\times 1+8}=\frac{2}{3}\]